# Xiao Liang Zhang Tiryaki Bae Shankar Yang Poon Dupoux Khudanpur Perera USA USA USA CASPER: A Large Scale Spontaneous Speech Dataset

### v1

@@ Abstract
The success of large language models has driven interest in developing similar speech processing capabilities. However, a key challenge is the scarcity of high-quality spontaneous speech data, as most existing datasets contain scripted dialogues. To address this, we present a novel pipeline for eliciting and recording natural dialogues and release our Stage 1 dataset with 200+ hours of spontaneous speech. Our approach fosters fluid, natural conversations while encouraging a diverse range of topics and interactive exchanges. Unlike traditional methods, it facilitates genuine interactions, providing a reproducible framework for future data collection. This paper introduces our dataset and methodology, laying the groundwork for addressing the shortage of spontaneous speech data. We plan to expand this dataset in future stages, offering a growing resource for the research community.
keywords:
Speech Processing, datasets, spontaneous speech, conversational speech1 Introduction
@@
To address this crucial gap, we introduce a comprehensive speech dataset named CAsual Speech in Peer Engagement Recordings, or CASPER, with 200 hours of recorded conversations, of which 158 hours are actual speech. Our goal is to expand this to 2,000 hours of natural, spontaneous dialogue, enriched with metadata such as speaker demographics and recording devices. By designing interactions that emulate typical telephone conversations, our dataset captures the spontaneity, turn-taking dynamics, and informal speech patterns essential for improving ASR performance in real-world conversational settings. The pipeline’s device-agnostic architecture allows for recording across various hardware setups, from professional audio equipment to common mobile devices, making it highly accessible for future data collection efforts. In this paper, we present our Stage 1 release, which comprises 200 hours of spontaneous speech along with partial transcriptions. Unlike existing datasets such as Switchboard [11] which primarily features shorter conversations and limited recording conditions, or GigaSpeech [10] which relies on existing media sources, our approach actively facilitates extended, natural dialogues that better capture the full complexity of human conversation. The longer conversation duration enables the development of more complex topics and natural interaction patterns, while our multi-device support ensures consistent quality across different recording environments. By releasing both the initial dataset and our flexible collection framework, we aim to establish a foundation for addressing the scarcity of spontaneous speech data, enabling researchers to study various aspects of natural conversation while contributing to this growing resource.
2 Related Work
2.1 Speech Large Language Model (Speech LLM)
@@
2.2 Existing Conversational Speech Corpora
@@
3 Data Collection
3.1 Data Collection Procedure
Our dataset comprises audio recordings of spontaneous conversations between pairs of participants, captured in real-world settings using their personal devices. To facilitate a structured yet natural data collection process, we developed a web-based platform where participants could initiate and record their conversations.111We open source our code at [link hidden for review].
Participants were recruited through emails and flyers, after which they registered on the platform, verified their email, and provided metadata, including age, native language, accent, and device type. Prior to participation, they reviewed and signed an informed consent form. A visualization of the pipeline is shown in Figure 1.
To encourage natural dialogue, participants selected conversation rooms labeled with broad topic descriptors (e.g., Favorite Activities). These prompts served as conversation starters rather than strict guidelines, allowing discussions to evolve freely. Each participant’s audio was recorded separately, preserving per-channel recordings for enhanced signal quality and post-processing flexibility. Upon completion, participants provided feedback through a survey assessing the ease and quality of the recording process.
@@ 3.1 Data Collection Procedure
3.2 Data Collection Format
@@
The conversations were recorded using the backend infrastructure of Daily.co222https://www.daily.co/. Each participant’s audio was recorded independently, preserving per-speaker channels to enable downstream processing tasks such as speaker diarization and analysis of overlapping speech. This per-channel recording approach enhances the dataset’s utility for speech processing research by providing cleaner, source-separated signals rather than a single mixed audio stream.
The recorded audio files were initially stored in an Amazon AWS S3 bucket in WebM format with a sample rate of 48 kHz. To ensure long-term accessibility and compatibility with standard speech processing tools, the files were subsequently transferred to an internal computing cluster and converted to WAV format at the same sample rate, with no loss in fidelity.
@@
User metadata was exported from the Firebase333https://firebase.google.com/ Realtime Database in JSON format. In order to protect sensitive information, personally identifiable data (e.g., names and geographic details) were stored separately, indexed by a unique User ID, and restricted to administrative access only. Before the dataset is made publicly available, appropriate de-identification procedures will be applied to ensure participant privacy.
3.3 Data Storage and Post-Processing
@@
To facilitate indexing and statistical analysis, the WAV recordings and their corresponding JSON metadata were integrated into a SQL database. Next, voice activity detection (VAD) and ASR were performed to estimate the ratio of speech to nonspeech for each recording, revealing that speech comprised about 78% of the audio.[21] The dataset was then refined using Label Studio444https://labelstud.io/—an open-source annotation platform—for human annotation, enabling fine-grained labeling of high-quality transcripts.
4 Data Overview
4.1 Metadata Analysis
@@
The dataset comprises a total of 208 participants, each contributing spontaneous speech recordings along with associated metadata, as shown in Table 1. The inclusion of demographic and technical details enhances the dataset’s utility, providing researchers with the necessary context to analyze speaker characteristics, recording conditions, and linguistic diversity. Below, we discuss the distribution of key metadata attributes in detail.
4.1.1 Speaker Demographics
@@
The participants exhibit a diverse range of linguistic backgrounds and regional accents. The majority (67.79%) reported speaking US English, reflecting the dataset’s primary demographic. However, a significant proportion of non-native and regionally influenced English varieties are also present, including Chinese Mandarin-influenced English (4.81%), UK English (5.29%), and Indian English (2.88%). Additionally, 14.42% of participants did not specify an accent, indicating either an omission or variability in self-identification.
@@ 4.1.2 Linguistic Diversity
4.1.3 Recording Conditions and Device Usage
4.1.4 Post-recording Survey Analysis
@@
4.2 Data Description
4.2.1 Overview
@@
4.2.2 Naturalness
As shown in Table 2, the transcript reflects key characteristics of spontaneous, natural conversations. Unlike scripted speech, natural dialogues often contain disfluencies such as fillers (”um”), restarts (”I was kind of, um, having, um…”), and self-corrections (”me, I’m not, I’m not someone that really travels a lot”). These elements indicate that speakers are thinking in real time, shaping their responses dynamically rather than delivering pre-formulated statements.
@@ 4.2.2 Naturalness
4.2.3 Long-Context Dependency
@@ 4.3 Benchmark
We computed word error rate (WER) using Whisper’s built-in English normalizer for both references and hypotheses. The results, shown in Table 3, indicate that Whisper-large-v3 achieves a WER of 0.31, while SeamlessM4T-large performs worse with a WER of 0.53. These relatively high error rates highlight the challenges of spontaneous, conversational speech, which includes disfluencies, informal phrasing, and long-context dependencies—characteristics that are underrepresented in the training data of these models. This suggests that our dataset is a valuable resource for improving ASR models, particularly for real-world conversational speech.
5 Conclusion
@@
In this paper, we introduced a new dataset of spontaneous conversational speech, totaling 200 hours of recorded audio, with 158 hours identified as actual speech. The dataset captures the natural flow of dialogue, including disfluencies, pauses, and speaker interactions, primarily between acquaintances. This makes it a valuable resource for developing more robust ASR systems capable of handling real-world conversational challenges.
@@
Evaluation of the baseline models on a randomly sampled subset of our dataset highlights the difficulties ASR systems face with spontaneous speech. This result underscores the need for datasets like ours to train models that can better handle natural conversations, disfluencies, and dynamic speech patterns. Our work provides a foundation for future research aimed at improving ASR systems, and we plan to expand the dataset and evaluate additional models to enhance performance further.

### v2

# CASPER: A Large Scale Spontaneous Speech Dataset

@@ Abstract
The success of large language models has driven interest in developing similar speech processing capabilities. However, a key challenge is the scarcity of high-quality spontaneous speech data, as most existing datasets contain scripted dialogues. To address this, we present a novel pipeline for eliciting and recording natural dialogues and release our dataset with 100+ hours of spontaneous speech. Our approach fosters fluid, natural conversations while encouraging a diverse range of topics and interactive exchanges. Unlike traditional methods, it facilitates genuine interactions, providing a reproducible framework for future data collection. This paper introduces our dataset and methodology, laying the groundwork for addressing the shortage of spontaneous speech data. We plan to expand this dataset in future stages, offering a growing resource for the research community.
Index Terms:
Speech Processing, datasets, spontaneous speech, conversational speechI Introduction
@@
To address this crucial gap, we introduce a comprehensive speech dataset named CAsual Speech in Peer Engagement Recordings, or CASPER, with 200 hours of recorded English conversations, of which 158 hours are actual speech. We collected a total of 200 hours of speech; however, this release contains the first 102 hours to support early research and development.111 https://huggingface.co/datasets/CASPER-SSSD/CASPER/ The remaining data will be released in conjunction with a future challenge to encourage research involvement and participation. This subset of the dataset is partitioned into train (80%) and test (20%) splits. To ensure the integrity of the evaluation, we release speaker metadata (e.g., accent, age) and labels for both splits. This is an initial effort to expand the dataset to 1,000 hours of natural, spontaneous dialogue, enriched with metadata such as speaker demographics and recording devices.
Central to this research is the investigation of interactions in typical telephone conversations. Our dataset captures the spontaneity, turn-taking dynamics, and informal speech patterns essential for improving downstream tasks such as Automatic Speech Recognition (ASR), speaker recognition, diarization, and topic detection performance in real-world conversational settings. Our data collection platform specifically targets pairs of individuals who already know each other, fostering natural, unforced conversations that can capture the properties of real-world conversations.
The pipeline’s device-agnostic architecture allows for recording across various hardware setups, from professional audio equipment to common mobile devices, making it highly accessible for future data collection efforts. In this paper, we present our initial release, which comprises a subset of 102 hours from a total of 200 hours of spontaneous speech, along with partial transcriptions. Unlike existing datasets such as Switchboard [11], which primarily features shorter conversations and limited recording conditions, or GigaSpeech [10], which relies on existing media sources, our approach actively facilitates extended, natural dialogues that better capture the full complexity of human conversation. The longer conversation duration enables the development of more complex topics and natural interaction patterns, while our multi-device support ensures consistent quality across different recording environments. By releasing both the initial dataset and our flexible collection framework, we aim to establish a foundation for addressing the scarcity of spontaneous speech data, enabling researchers to study various aspects of natural conversation while contributing to this growing resource.
II Related Work
II-A Speech Large Language Model (Speech LLM)
@@
II-B Existing Conversational Speech Corpora
@@
Also being part of the Scalable Spontaneous Speech Dataset (SSSD) project, [21] adopts a crowdsourced approach to spontaneous speech collection. Like our methodology, they utilize a web-based platform for data collection, ensuring accessibility and ease of use. While their pipeline relies on Amazon Mechanical Turk (MTurk) for participant recruitment, often involving strangers, whereas our approach avoids crowdsourcing marketplaces in favor of more targeted recruitment, specifically acquaintances. This allows us to foster more natural interactions, reducing the risk of rushed or disengaged contributions and resulting in more meaningful and coherent dialogues.
III Data Collection
III-A Data Collection Procedure
Our dataset comprises audio recordings of spontaneous conversations between pairs of participants, captured in real-world settings using their personal devices. To facilitate a structured yet natural data collection process, we developed a web-based platform where participants could initiate and record their conversations.222We open-source our code at https://anonymous.4open.science/r/sssd-D5CA/README.md.
Participants were recruited through emails and flyers, after which they registered on the platform, verified their email, and provided metadata, including age, native language, accent, and device type. Prior to participation, they reviewed and signed an informed consent form. A visualization of the pipeline is shown in Fig. 1. Pairs of peer participants already familiar with each other were targeted in the recruitment process to facilitate natural conversations. To prioritize the study of natural, spontaneous conversations, our dataset places secondary emphasis on diversity in gender, language, and other demographic factors.
Each participant was able to schedule recordings per their own timeline using the developed scheduling tool seen in Fig. 2.
To encourage natural dialogue, participants selected conversation rooms labeled with broad topic descriptors (e.g., Favorite Activities) seen in Fig. 3.
These prompts served as conversation starters rather than strict guidelines, allowing discussions to evolve freely. Each participant’s audio was recorded in a separate channel for enhanced signal quality and post-processing flexibility. Upon completion, participants provided feedback through a survey assessing the ease and quality of the recording process. The recording screen from the participants’ end and the post-recording survey can be seen in Fig. 5 and Fig. 5 respectively.
@@ III-A Data Collection Procedure
III-B Data Collection Format
@@
The conversations were recorded using the backend infrastructure of Daily.co333https://www.daily.co/. Each participant’s audio, recorded in a separate channel, enables downstream processing tasks such as automatic speech recognition per-speaker, speaker diarization and analysis of overlapping speech. This per-channel recording approach enhances the dataset’s utility for speech processing research by providing cleaner, source-separated signals rather than a single mixed audio stream. The recorded audio files were initially stored in an Amazon AWS S3 bucket in WebM format with a sample rate of 48 kHz. To ensure long-term accessibility and compatibility with standard speech processing tools, the files were subsequently transferred to an internal computing cluster and converted to WAV format at the same sample rate, with no loss in fidelity.
@@
User metadata was exported from the Firebase444https://firebase.google.com/ Realtime Database in JSON format. In order to protect sensitive information, personally identifiable data (e.g., names and geographic details) were stored separately, indexed by a unique User ID, and restricted to administrative access only.
III-C Data Storage and Post-Processing
@@
To facilitate indexing and statistical analysis, the WAV recordings and their corresponding JSON metadata were integrated into a SQL database. Next, voice activity detection (VAD) using Silero-v4 [22] and ASR using Whisper-large-v3 [23] were performed to estimate the ratio of speech to nonspeech for each recording, revealing that speech comprised about 78% of the audio. The dataset was then refined using Label Studio555https://labelstud.io/—an open-source annotation platform—for human annotation, enabling fine-grained labeling of high-quality transcripts.
III-D De-Identification
To protect the privacy of the participants the audio files, as well as the metadata, were de-identified. To achieve this, each recording was transcribed along with word-level timestamps using WhisperX [24]. After this, each transcription was searched to find any occurrence of the participant provided metadata. The identified metadata information, such as names and surnames, were then replaced with silences in the recordings based on their word-level timestamps.
IV Data Overview
IV-A Metadata Analysis
@@
The dataset comprises a total of 208 participants, each contributing spontaneous speech recordings along with associated metadata, as shown in Table I. The inclusion of demographic and technical details enhances the dataset’s utility, providing researchers with the necessary context to analyze speaker characteristics, recording conditions, and linguistic diversity. Below, we discuss the distribution of key metadata attributes in detail.
IV-A1 Speaker Demographics
@@
The participants exhibit a diverse range of linguistic backgrounds and regional accents. The majority (67.79%) reported speaking US English, reflecting the dataset’s primary demographic. However, a significant proportion of non-native and regionally influenced English varieties are also present, including Chinese Mandarin-influenced English (4.81%), UK English (5.29%), and Indian English (2.88%). Additionally, 14.42% of participants did not specify an accent, indicating either an omission or variability in self-identification. The participants’ accent and native language are based on their self-identification, for example, the number of speakers with an Arabic accent may differ from the number with Arabic as their native language.
@@ IV-A2 Linguistic Diversity
IV-A3 Recording Conditions and Device Usage
IV-A4 Post-recording Survey Analysis
IV-B Data Description
IV-B1 Overview
@@
IV-B2 Naturalness
As shown in Table II, the transcript reflects key characteristics of spontaneous, natural conversations. Unlike scripted speech, natural dialogues often contain disfluencies such as fillers (”um”), restarts (”I was kind of, um, having, um…”), and self-corrections (”me, I’m not, I’m not someone that really travels a lot”). These elements indicate that speakers are thinking in real time, shaping their responses dynamically rather than delivering pre-formulated statements.
@@ IV-B2 Naturalness
IV-B3 Long-Context Dependency
@@ IV-C1 Automatic Speech Recognition
We computed word error rate (WER) using Whisper’s built-in English normalizer for both references and hypotheses. The results, shown in Table III, indicate that Whisper-large-v3 achieves a WER of 0.31, while SeamlessM4T-large performs worse with a WER of 0.53. These relatively high error rates highlight the challenges of spontaneous, conversational speech, which includes disfluencies, informal phrasing, and long-context dependencies—characteristics that are underrepresented in the training data of these models. This suggests that our dataset is a valuable resource for improving ASR models, particularly for real-world conversational speech.
IV-C2 Diarization
We evaluated the 102 hours of publicly available conversations by 157 distinct speakers on the diarization task. Each conversation consists of two time-aligned audio files, one from each participant’s end. We combined the two files into a single-channel recording, which was then used for diarization evaluation.
WhisperX [24] is an ASR model built on Whisper that employs forced alignment with wav2vec2 [26] to generate word-level timestamps. To create the ground truth, speech segments were identified using WhisperX-generated timestamps on each participant’s recording independently, for every conversation. These timestamps were mapped onto the combined conversation to be used as reference labels.
For the benchmark of diarization performance, we used two state-of-the-art diarization models: pyannote.audio-v3.1[27, 28] and the NeMo Cascaded Speaker Diarization System[29]. Pyannote-audio v3.1 is a fully automatic diarization system that separates speakers without relying on an external VAD model, and is widely adopted in diarization tasks and used for benchmarking. The model uses a combination of EEND [30] diarization and unsupervised clustering.
The NeMo framework, developed by NVIDIA, includes a neural cascaded diarization architecture combining specialized components for each sub-task. Our benchmark uses TitaNet-Large[31], a speaker embedding model trained on thousands of hours of speech, and MarbleNet[32], a VAD model trained on over thousands of hours of multi-language speech data. Together, these components enable robust speaker segmentation and clustering, making the NeMo Neural Diarizer well suited for evaluating this dataset.
We measure Diarization Error Rate (DER) and Jaccard Error Rate (JER) computed without a forgiveness collar and with regions of overlapped-speech included in the evaluation. The results metrics can be seen in Table IV as well as the histogram of error rates across recordings in Fig. 6. The histogram shows the mean DER and JER across bins of recording durations, enabling the evaluation of diarization performance that is related to length of the recording. The means are over each bin only, the number of files in each bin affect the overall error rates seen in Table IV.
| Model | Metric | Mean (%) | Median (%) | 5% / 95% (%) |
| Pyannote | DER | 25.93 | 23.24 | 16.80 / 51.61 |
| JER | 29.63 | 24.59 | 15.48 / 57.20 | |
| NeMo | DER | 30.77 | 28.23 | 16.80 / 53.48 |
| JER | 33.87 | 30.85 | 17.62 / 60.38 |
The relatively high error rates on our dataset suggest an increased difficulty of diarization on conversational speech with disfluencies, overlaps, and frequent turn-taking dynamics that are underrepresented in existing resources. The natural setting of recordings, where each participant records from their own device in their own environment, also introduces new obstacles for the system. These results indicate that our dataset presents real-world challenges in diarization systems and can be a suitable and valuable resource in the speech community to improve the accuracy and robustness of diarization models for conversational speech.
V Conclusion
@@
In this paper, we introduced a new dataset of spontaneous conversational speech, totaling 200 hours of recorded audio, with 158 hours identified as actual speech. The dataset captures the natural flow of dialogue, including disfluencies, pauses, and speaker interactions, primarily between acquaintances. This makes it a valuable resource for developing more robust downstream task systems (ASR, diarization, topic detection, among others) capable of handling real-world conversational scenarios. This first release comprises a subset of 102 hours of audio, publicly available.
@@
Evaluation of the baseline models on a randomly sampled subset of our dataset highlights the difficulties ASR and diarization systems face with spontaneous speech. This result underscores the need for datasets like ours to train models that can better handle natural conversations, disfluencies, and dynamic speech patterns. Our work provides a foundation for future research aimed at improving downstream task systems, and we plan to expand the dataset and evaluate additional models to enhance performance further.